# AnyBipe: An End-to-End Framework for Training and Deploying Bipedal Robots Guided by Large Language Models

### v1

@@ IV-A Module Analysis
| Standing still1 | -50.35 | -6.16 | - 11.20 | 5.8 (30) s |
| Survival time | 0.86 | 0.30 | 0.30 | 30 (30) s |
@@
IV-B Framework Analysis
The experimental setup involved first establishing a baseline using a manually designed reward function. We then trained our model from scratch in flat and complex terrain environments, sed to simulate the general situation of users first using the AnyBipe framework. The robot was trained to track speed commands, avoid falls, and walk on various terrains using a simplified reward function with basic components like velocity tracking, action rate, and balance judgment, without scale-related information. Over five rounds ( samples each), the best model from each round was tested in Gazebo and real-world scenarios. The basic training framework occupies 7G of GPU memory, and the corresponding training time is 79 hours. As shown in Fig. 8, initial iterations had issues like excessive movement (Iteration 0) and unnatural joint postures (Iterations 1 and 2), but these were corrected by Iteration 5, achieving a natural gait.
To demonstrate the effectiveness of the LLM’s reward function improvements during training, we evaluate its performance using reward success and terrain level in tasks like velocity tracking, safe walking, and navigating complex terrains. After comparing five iterations with manually designed reward functions in Fig. 9, we found that the LLM’s reward function outperformed the manual version after just two iterations. Each subsequent iteration further enhanced training speed and performance, all achieved without human intervention.
To verify that the final trained policy not only performs well in the lab but also adapts to real-world terrains, we conducted walking tests across five different surfaces: the experimental site, carpet, hard ground, grass, and stairs. The experiments demonstrated that the model trained and deployed by AnyBipe possesses the ability to walk on various terrains. In contrast, the manually designed reward function used in the experiment failed to achieve Sim-to-real transfer, indicating that AnyBipe can independently resolve the Sim-to-real problem.
@@ V Conclusion
AnyBipe proposes an end-to-end framework for training and deploying bipedal robots, which utilizes a state-of-the-art LLM to design reward functions for specific tasks. The framework provides interfaces that allow users to supply coding references and integrate pre-existing models to assist in training. Additionally, it incorporates feedback from both simulated and real-world test results, enabling the execution of Sim-to-Real tasks entirely without human supervision. We validated the effectiveness of each module, as well as the system’s ability to guide the robot in learning locomotion in complex environments, continuously improving the model by either designing new reward functions from scratch or refining existing ones. Furthermore, this framework exhibits potential for transferability to other robotic task planning scenarios. Our future work will focus on improving the current framework in three key areas: first, extending it to a broader range of robotic applications to verify its generalizability; second, testing its effectiveness across more tasks beyond locomotion; and third, enhancing the model evaluation process by incorporating visual and conceptional feedback to achieve more comprehensive state estimation.
APPENDIX
V-A Proof of section 3.2
We have KL divergence between normal distribution with PDF , and Dirac distribution with PDF , the KL divergence can be written as
We first prove that . Let , , where is the step function, we have
Therefore we can approximately say that
Then we show why we don’t calculate the normal KL divergence . This is because the term
is , there is no way for normal calculation. Since the backward propagation functions normally, we choose the inverse form, which approximately measures the difference between two distributions.
Also, in practical, we tend to ignore the term which controls the size of , using this term directly instead:

### v2

@@ IV-B Module Analysis
Standing still1 | -5.35 | -6.16 | - 11.20 | 5.8 (30) s |
Survival time | 0.86 | 1.0 | 1.0 | 30 (30) s |
@@
For the safety check criterion , we define two levels: the base criterion, which includes falling off, self-collision (not implemented for P1222Contact detection extraction is implemented in MuJoCo only), and violating DoF and torque limits; and the warning criterion, which includes high feet contact forces (not implemented for P1), high action rate, high total torque, and large gravity projection deviation. Policies passing the base safety tests are considered deployable, and those without warnings are regarded as safe. We conducted 3 training groups for each robot, generating 240 samples per robot ( results per run). Models passing the base test are labeled as positive, and those passing both tests as strict positive. The ground truth is established by manually evaluating each policy’s behavior in simulation to identify true positives and negatives. We then compute the average precision (AP) and average recall (AR) to assess performance. We also list another success rate (SR) defined as the percentage of policy passed the safety check. Evaluation proved that properly implemented can identify whether policy is safe under most scenarios. This also shows that prompts with better initial prompts leads to higher success rate.
Name | AP | AR | S-AP | S-AR | SR (S-SR) |
P1 Pointfoot | 94.3 | 96.4 | 100.0 | 92.6 | 58.7 (26.2) |
Unitree H1 | 100.0 | 97.5 | 93.3 | 84.0 | 65.4 (18.8) |
Turin Robot | 97.6 | 96.1 | 83.8 | 93.9 | 52.9 (15.4) |
IV-C Framework Analysis
The experimental setup is summarized in Table V. For each type of experiment, we present results using different LLM bases333The choice of LLMs varies for different robots. P1 showed promising results with GPT-4o, and alternative models did not significantly improve performance. We intended to use Deepseek-R1, but its official API was unstable during the experiment (Jan. 2025 to Feb. 2025), so we opted for Claude-3.5-Sonnet instead.. Detailed results and setup for similar experiments can be found on GitHub.
Name | LLM | Train Iter | Sim Env | Real Env |
P1 Pointfoot | GPT-4o | 5,000 | Gazebo | |
Unitree H1 | Claude-3.5 | 2,000 | MuJoCo | |
Turin Robot | Deepseek-R1 | 2,000 | MuJoCo |
Name | Iter 0 | Iter 1 | Iter 2 | Iter 3 | Iter 4 | Human Engineered |
P1 Pointfoot (iter=5000) | 14.7 (5.5) | 7.1 (6.3) | 9.2 (8.7) | 15.5 (11.2) | 25.6 (20.8) | 19.2 (19.2) |
Unitree H1 (iter=2000) | 0.0097 (0.0035) | 21.2 (12.4) | 25.2 (21.7) | 25.0 (22.4) | 25.8 (24.9) | 23.7 (23.7) |
Turin Robot (iter=2000) | 10.2 (1.96) | 52.3 (42.9) | 58.1 (51.8) | 66.3 (55.4) | 67.0 (57.8) | None (None) |
Fig. 5 describe the best training result iterating over framework iterations, along with feedback indicating different levels of safety violation (red for not safe, yellow for potential problem warning, and green for passed check). P1 passes all check and all reality deployments are listed. The other two are listed in training-simulation pairs. These experiments prove that AnyBipe is capable of guiding LLM to realize what problems might occur in certain reward function implementation, and can act correspondingly to solve the problem. By iterations the unsafe problems are properly handled while maintaining code set effective.
We visualize the reward and terrain level curves for P1 training, comparing them to human-engineered reward curves extracted from TensorBoard logs in Fig. 8. For the three experiments, we report the average success rate at the midpoint and end of training for 5 iterations, comparing them with human-engineered metrics (considered state-of-the-art, as they are available open-source) in Table VI. Our results demonstrate that AnyBipe can identify more effective reward function combinations after several iterations without human intervention. The framework can also autonomously explore suitable code implementations in zero-shot scenarios when no reference is provided (as shown in the Turin task). This highlights the framework’s potential for training newly designed robot configurations.
Experiments in Fig. 8 indicates that the trained policy can be truly deployed in reality, and maneuvers different kinds of terrain types.
@@ V Conclusion
AnyBipe presents an end-to-end framework for training and deploying bipedal robots, leveraging state-of-the-art Large Language Models (LLMs) to design reward functions tailored to specific tasks. The framework provides interfaces that allow users to supply coding references and integrate pre-existing models to facilitate the training process. Furthermore, it incorporates feedback from both simulated and real-world test results, enabling the execution of Sim-to-Real tasks without human supervision while also offering improvement directions to the LLMs. We have validated the effectiveness of each module, as well as the system’s capacity to guide the robot in learning locomotion in complex environments, progressively improving the model by either creating new reward functions from scratch or refining existing ones. Moreover, this framework demonstrates potential for transferability to other robotic task planning scenarios.
Despite these strengths, there remain areas for further improvement. Treating observations and reward terms solely as human factors may not fully capture human preferences, and visual factors should be considered as well. Also, the framework automates the whole training process, but still requires human to implement the basic RL environment following legged gym and our template, not generating environment directly from LLM. Additionally, incorporating Vision-Language Models (VLMs) into the safety criterion could improve the precision of judgment.
Future work will focus on advancing the framework in key areas: first, expanding its application to a broader range of robotic tasks to assess its generalizability, testing its effectiveness beyond locomotion; second, introducing a supervisory process during the training stage to allow the framework to autonomously determine the duration of reinforcement learning training, rather than relying on a fixed total length; and third, refining the model evaluation process by incorporating visual and conceptual feedback to achieve a more comprehensive state estimation.